A generic solid, X, has a molar mass of 61.6 g/mol. In a constant-pressure calorimeter, 18.3 g of X is dissolved in 383 g of water at 23.00 °C.X(s) ͢ X(aq)The temperature of the resulting solution rises to 29.10 °C. Assume the solution has the same specific heat as water, 4.184 J/(g·°C), and that there's negligible heat loss to the surroundings. How much heat was absorbed by the solution?What is the enthalpy of the reaction?
Answer #1

2 gic HEAT Rolenseo, q = MCAT mass of solution, m= 18.3 g+ 383g = 40103 g Temperature Ctange, at = (29.10 – 23.00) = 6-10°c •